The Virial Equation of State can be used to descri

What are the units of Z, B, C, and D?
Answer #1

The virtual equation of state is: pVm/RT = 1 + B/Vm + C/Vm2 + D/Vm3 + .......

where B,C,D are functions of temperature. So the units of B,C,D will be oC or oF

In the given equation Z = pV/RT. The general equation is pV = nRT i.e n= pV/RT where p is pressure,V is volume,R is gas constant, T is temperature and n is number of moles. In the given equation, n is represented as Z

Therefore, unit of Z is mole.